OpenStudy (anonymous):

simplfy if possible: (x^2+9x+20/4x) / (5x+25/2x^2)

OpenStudy (anonymous):

Have you multiplied by the reciprocal yet?

OpenStudy (anonymous):

no not yet

OpenStudy (anonymous):

i confused myself...so i need to know what to do after i do the reciprocal

OpenStudy (anonymous):

Factor the x^2+9x+20 out. Remember how to do this?

OpenStudy (anonymous):

yes i remeber that and i came up with (x+5) (x+4)

OpenStudy (anonymous):

4 times 5 is 20, 4 plus 5 is 9. You get (x+4)(x+5)

OpenStudy (anonymous):

All right, and then on the denominator is 5x+25. What can you do with that to cancel out something from (x+4)(x+5)?

OpenStudy (anonymous):

\[\frac{\frac{x^2+9x+20}{4x}}{\frac{5x+25}{2x^2}}\] \[\frac{x^2+9x+20}{4x}*\frac{2x^2}{5x+25}\] \[\frac{(x+4)(x+5)}{4x}*\frac{2x^2}{5(x+5)}\] \[\frac{x(x+4)}{2*5}\] \[\frac{x^2+4x}{10}\]

OpenStudy (anonymous):

That must have been a pain to type

OpenStudy (anonymous):

thanks everyone for the help!!

OpenStudy (anonymous):

np